OpenStudy (anonymous):

In an isosceles triangle, the base is 7 more than one-half times one of the legs. If the perimeter is 22 cm, find the length of each leg.

OpenStudy (anonymous):

You get an equation of the form: L + L + (L/2 + 7) = 22 -> 5L/2 + 7 = 22. Can you solve this for L?

OpenStudy (anonymous):

l=6

OpenStudy (anonymous):

Excellent! Good job!

OpenStudy (anonymous):

now how do you show how to get the base

OpenStudy (anonymous):

ok, since the two legs are each "L", the base is 7 more than half of a leg. So, you take a leg, "L", and first make it half, so you get L/2, but you're not done yet, because the base is 7 more than this, so L/2 + 7 for the base.

OpenStudy (anonymous):

ok thanks so much for your help

OpenStudy (anonymous):

You are quite welcome!

OpenStudy (anonymous):

how did you get 5L?

OpenStudy (anonymous):

tcarrol010 got not really 5L but (5/2)L from adding L+L+(1/2)L: (1+1+1/2)L = (5/2)L. Do u now see where tcarroll010 got this?

OpenStudy (anonymous):

Still not sure

OpenStudy (anonymous):

What are you not sure about? is it how he/she got the equation? or is it still how he/she got (5/2)L?

OpenStudy (anonymous):

Here is a diagram. Does this help? 

OpenStudy (anonymous):

In an isosceles triangle the 2 legs are always equal.